# iConFormer: Dynamic Parameter-Efficient Tuning with Input-Conditioned Adaptation

### v1

@@ Abstract
Transfer learning based on full fine-tuning (FFT) of the pre-trained encoder and task-specific decoder becomes increasingly complex as deep models grow exponentially. Parameter efficient fine-tuning (PEFT) approaches using adapters consisting of small learnable layers have emerged as an alternative to FFT, achieving comparable performance while maintaining high training efficiency. However, the inflexibility of the adapter with respect to input instances limits its capability of learning task-specific information in diverse downstream tasks. In this paper, we propose a novel PEFT approach, input-Conditioned transFormer, termed iConFormer, that leverages a dynamic adapter conditioned on the input instances. To secure flexible learning ability on input instances in various downstream tasks, we introduce an input-Conditioned Network (iCoN) in the dynamic adapter that enables instance-level feature transformation. To be specific, iCoN generates channel-wise convolutional kernels for each feature and transform it using adaptive convolution process to effectively capture task-specific and fine-grained details tailor to downstream tasks. Experimental results demonstrate that by tuning just 1.6% to 2.8% of the Transformer backbone parameters, iConFormer achieves performance comparable to FFT in monocular depth estimation and semantic segmentation, while outperforming it in image classification and instance segmentation. Also, the proposed method consistently outperforms recent PEFT methods for all the tasks mentioned above.
@@
Figure 1: Quantitative comparison with full fine-tuning (FFT) and PEFT approaches. The top graph compares depth prediction errors on NYU-v2 (Silberman et al. 2012), while the bottom graph presents performance for semantic segmentation on ADE20K (Zhou et al. 2017) and instance segmentation on COCO (Lin et al. 2014). iConFormer consistently outperforms recent PEFT methods in all dense tasks and surpasses FFT in instance segmentation.
Introduction
@@
As deep neural networks (DNNs) grow increasingly complex, transfer learning—fine-tuning pre-trained models with task-specific data for downstream tasks—has become a widely adopted solution across diverse applications, including image classification, semantic segmentation, and object detection, to name a few. For instance, the model consisting of the pre-trained encoder (Xie et al. 2022b; He et al. 2022a) and task-specific decoder is fine-tuned, achieving remarkable performance gain when compared to the model trained from scratch. However, training large, complex models separately for each task and dataset is highly inefficient. Recently, parameter efficient fine-tuning (PEFT) approaches (Zaken, Ravfogel, and Goldberg 2021; Qi et al. 2022; Hu et al. 2021) that maximize the efficiency in terms of training parameters have emerged as an alternative to the above-mentioned full fine tuning (FFT) methodologies, achieving competitive performance even with limited computing resources while simplifying the training processes and deployment.
@@
This remarkable progress in vision tasks is primarily driven by approaches including prompt tuning (Lester, Al-Rfou, and Constant 2021) and adapter (Houlsby et al. 2019), which have been successfully applied in natural language processing (NLP) tasks. Visual prompt tuning (Jia et al. 2022) is the first study to explore the potential of prompt tuning in visual recognition tasks, laying the foundation for the prompt tuning in the field of computer vision. In addition, the adapter-based PEFT methods (Chen et al. 2022; Yin et al. 2023) achieve significant training efficiency by applying the adapter used in NLP to the Vision Transformer (ViT) and its variants (Dosovitskiy et al. 2021; Liu et al. 2021, 2022).
@@
While most PEFT-based approaches yield comparable performance to baseline methods using the FFT in the image classification task, they do not yet provide sufficiently satisfactory performance to compete with the FFT in other complex downstream tasks. Additionally, the scalability of prompt-based methods (Jia et al. 2022; Bahng et al. 2022) is significantly limited, leading to considerable performance degradation as the number of learnable parameters increases, as reported in (Chen et al. 2022). Adapter-based approaches (Chen et al. 2022; Hu et al. 2021; Yin et al. 2023) incorporate lightweight modules to reduce the number of trainable parameters while maintaining stable performance across a range of trainable parameter scales. However, the adapters always apply the same transformation to input features, ignoring individual characteristics of input instances. This may not be an issue when fully tuning DNNs, but it could be a limiting factor in improving performance in adapter-based PEFT methods. Namely, the inflexibility with respect to input instances exacerbates the transfer capability of adapter-based models with small learnable parts to downstream tasks, limiting their ability to capture unique and task-specific information.
@@
Additionally, the Vision Transformer (ViT) backbone (Dosovitskiy et al. 2021) used in adapter-based models tends to focus on global information rather than fine-grained local details within an image. While this limitation can be partially addressed by employing the Swin Transformer (Liu et al. 2021), which utilizes local attention mechanisms, the constraint on the number of learnable parameters in PEFT still restricts the Swin Transformer’s capability to effectively capture local features. Consequently, this negatively affects the performance in dense prediction tasks that require fine-grained information.
@@
To address the aforementioned issues, we propose a novel PEFT approach, input-Conditioned transFormer, termed iConFormer, that leverages a dynamic adapter where parameters are adjusted input instance-level. Unlike recent adapter-based approaches (Chen et al. 2022; Hu et al. 2021; Yin et al. 2023), iConFormer introduces an input-Conditioned Network (iCoN) that dynamically generates the parameters for each input feature in the adapter. This approach enables for effectively capturing task-specific and local details while keeping the number of learnable parameters small. The effectiveness of our method is evidenced by the quantitative analysis in Figure 1, which shows that iConFormer captures task-specific information more effectively than conventional methods. We also analyze the capability to capture fine-grained details by visualizing attention maps in Figure 3.
iConFormer achieves performance competitive to full fine-tuning (FFT) on both classification and dense prediction tasks including monocular depth estimation, semantic segmentation, and instance segmentation with only additional 1.6% to 2.8% backbone parameters. Our method even surpasses FFT for the image classification on CIFAR100 (Krizhevsky, Hinton et al. 2009) and the challenging instance segmentation task on COCO (Lin et al. 2014). Additionally, iConFormer also outperforms conventional PEFT methods for monocular depth estimation task on NYU-v2 (Silberman et al. 2012), demonstrating the effective utilization of pre-trained backbone parameters with additional learnable parameters dynamically finetuned for specific tasks. In summary, our contributions are threefold:
@@
Figure 2: Comparison of Full Fine-Tuning (FFT) and the proposed Parameter Efficient Fine-Tuning (PEFT) using iConFormer. (a) FFT, where all parameters are updated during training. (b) Our PEFT (iConFormer), where adapters sequentially are attached after the MLP layer in the Transformer. Inside the adapters, an Input-Conditioned Network (iCoN) generates input-conditioned convolutional kernels in a channel basis. By convolving features with these kernels, iCoN adaptively refines the features in accordance with the specific properties of the input, thereby enhancing the model’s capability to effectively process diverse data in the downstream tasks.
Related Work
Transformer in Vision.
@@
Transformers, initially designed for Natural Language Processing (NLP) tasks such as machine translation (Vaswani 2017) and text generation (Devlin 2018), have achieved significant success in these areas. This success has led to a shift towards computer vision, starting with the Vision Transformer (ViT) (Dosovitskiy et al. 2021). Subsequently, various Transformer-based models (Liu et al. 2021; Xie et al. 2021; Carion et al. 2020; Lee et al. 2022, 2023; Ranftl, Bochkovskiy, and Koltun 2021; Hong et al. 2022) have achieved notable advancements in tasks including image classification (Krizhevsky, Sutskever, and Hinton 2012), semantic segmentation (Long, Shelhamer, and Darrell 2015; Chen et al. 2017; Cho et al. 2024; Kim, Choi, and Min 2022), object detection (Girshick et al. 2014; Ren et al. 2016; Redmon et al. 2016), image restoration (Dong et al. 2015; Liu, Breuel, and Kautz 2017), and depth estimation (Godard et al. 2019; Choi et al. 2021). Furthermore, transformers have significantly advanced vision recognition through large-scale pretraining (Chen, Xie, and He 2021; Caron et al. 2021; He et al. 2022a). However, their larger size compared to previously prevalent CNN backbones presents challenges for fine-tuning on specific tasks. In this context, our work explores methods to adapt pre-trained transformers into target tasks in a more effective and efficient way.
Parameter Efficient Fine Tuning.
@@
Parameter Efficient Fine-Tuning (PEFT) methods enable the adaptation of large self-supervised pre-trained models (Xie et al. 2022c; He et al. 2022b; Yi et al. 2022; Choi et al. 2024a, b) to specific tasks without the need to train the entire model. In Natural Language Processing (NLP), notable approaches include adapter methods (Houlsby et al. 2019), which integrate small modules into the model while keeping the pre-trained parameters frozen, with only the added modules being fine-tuned. Additionally, other methods involve tuning specific components, such as bias or normalization layers (Zaken, Ravfogel, and Goldberg 2021; Qi et al. 2022), utilizing learnable prompt tokens (Lester, Al-Rfou, and Constant 2021), or applying low-rank matrix approximations (Hu et al. 2021) to update parameters more efficiently.
@@
In computer vision, PEFT techniques inspired by NLP have shown significant progress. For instance, VPT (Jia et al. 2022) is the first method to apply prompt tuning approaches to visual classification tasks. AdaptFormer (Chen et al. 2022) employs a parallel adapter framework to enhance the effectiveness of parameter efficient fine-tuning for visual classification. KAdaptation (He et al. 2023) optimizes the adapter using Kronecker products. In addition, LoRand (Yin et al. 2023) employs multi-branch low-rank adapters to achieve impressive performance on dense prediction tasks. Our work focuses on adapter methods, proposing new approaches that aim to achieve performance comparable to or exceeding that of full fine-tuning across various downstream tasks.
Preliminary
Vision Transformer and its Variants
@@
Vision Transformer (ViT) (Dosovitskiy et al. 2021), modified from the Transformer (Vaswani 2017) proposed in NLP, integrates image patches and positional encodings to capture spatial information. It consists of a patch embedding layer and multiple sequential encoder blocks, as depicted in Fig. 2(a). Given a batch of images $x\in\mathbb{R}^{B\times H\times W\times 3}$, the patch embedding layer transforms $x$ into sequential patches $x_{p}\in\mathbb{R}^{B\times M\times(P^{2}C)}$, where $H\times W$ is an image resolution, and $B$ is a batch size. $P\times P$ is the resolution of an image patch, $C$ is the output channel, and $M=HW/P^{2}$ is the number of image patches. The patches are linearly embedded into $D$ dimensions to generate the final input $x_{in}\in\mathbb{R}^{B\times M\times D}$.
@@
In the Transformer encoder block, $x_{in}$ is first normalized using LayerNorm (Ba, Kiros, and Hinton 2016), and then processed by a multi-head self-attention layer (MHSA). The output is combined with $x_{in}$ via a residual connection:
@@
This process is repeated $N$ times in the encoder block. In ViT, the self-attention mechanism captures global features by evaluating relationships between all image patches, enabling a comprehensive understanding of complex dependencies. Advancing this approach, Swin Transformer (Liu et al. 2021, 2022) introduces hierarchical attention with shifted windows, which enhances both computational efficiency and feature representation. Other variants (Wang et al. 2021; Xie et al. 2021; Li et al. 2023; Xia et al. 2024) leverage multi-scale feature extraction for specific vision tasks, demonstrating the adaptability of Transformer models to diverse computer vision challenges.
PEFT Methods
@@
Parameter efficient fine-tuning (PEFT) methods, such as prompt tuning (Lester, Al-Rfou, and Constant 2021; Jia et al. 2022), low-rank adaptation (Hu et al. 2021), and adapters (Houlsby et al. 2019; Chen et al. 2022), are designed to reduce the number of trainable parameters needed for fine-tuning large models. Here, we briefly review adapter-based PEFT methods related to our approach, which will be detailed in the following section. Adapters introduce small, trainable modules between the layers of the pre-trained model. For instance, if the original model’s transformation is defined as $h=W\cdot x$, adapters modify this operation to $h=W\cdot x+A\cdot g(B\cdot x)$, where $B$ represents the down-projection of the input features, $A$ indicates the up-projection back to the original space, and $g(\cdot)$ denotes a non-linearity. These approaches allow for effective task adaptation while minimizing the number of parameters that need to be updated.
Figure 3: Comparison of Attention Maps from AdaptFormer and iConFormer. We visualize the attention maps using attention rollout (Abnar and Zuidema 2020). The top row represents input images, and the middle and bottom rows present the attention maps generated by AdaptFormer (Chen et al. 2022) and iConFormer, respectively, with both using Swin Transformer backbone (Liu et al. 2021). iConFormer more accurately delineates object regions and captures fine-grained semantics, compared to the AdaptFormer.
Proposed Method
@@
Conventional adapter-based methods (Houlsby et al. 2019; Chen et al. 2022) rely on static parameter tunings, which may constrain their ability to capture diverse input features with high precision. To address these limitations, we propose iConFormer, a novel framework incorporating the input-Conditioned Network (iCoN), as illustrated in Figure 2 (b). Unlike conventional adapters, iConFormer dynamically generates convolutional kernels tailored to the unique characteristics of each input. This dynamic generation facilitates more accurate and flexible feature extraction, introducing locality inductive biases into the pre-trained Transformer encoder. This approach enables the model to effectively capture both global and local features within a parameter-efficient framework. Consequently, iConFormer not only improves the model’s performance but also ensures a more precise adaptation to varying input conditions.
Input-Conditioned Network (iCoN)

### v2

@@ Abstract
Transfer learning based on full fine-tuning (FFT) of the pre-trained encoder and task-specific decoder becomes increasingly complex as deep models grow exponentially. Parameter efficient fine-tuning (PEFT) approaches using adapters consisting of small learnable layers have emerged as an alternative to FFT, achieving comparable performance while maintaining high training efficiency. However, the inflexibility of the adapter with respect to input instances limits its capability of learning task-specific information in diverse downstream tasks. In this paper, we propose a novel PEFT approach, input-Conditioned transFormer, termed iConFormer, that leverages a dynamic adapter conditioned on the input instances. To secure flexible learning ability on input instances in various downstream tasks, we introduce an input-Conditioned Network (iCoN) in the dynamic adapter that enables instance-level feature transformation. To be specific, iCoN generates channel-wise convolutional kernels for each feature and transform it using adaptive convolution process to effectively capture task-specific details tailored to downstream tasks. Experimental results demonstrate that by tuning just 1.6% to 2.8% of the Transformer backbone parameters, iConFormer achieves a performance comparable to FFT in monocular depth estimation and semantic segmentation, while outperforming it in image classification and instance segmentation. Additionally, the proposed method consistently outperforms recent PEFT methods for all the tasks mentioned above.
@@
Figure 1: Quantitative comparison with full fine-tuning (FFT) and PEFT approaches. The top graph compares depth prediction errors on NYU-v2 [55], while the bottom graph presents performance for semantic segmentation on ADE20K [68] and instance segmentation on COCO [45]. iConFormer consistently outperforms recent PEFT methods in all dense tasks and surpasses FFT in instance segmentation.
1 Introduction
@@
As deep neural networks (DNNs) grow increasingly complex, transfer learning—fine-tuning pre-trained models with task-specific data for downstream tasks—has become a widely adopted solution across diverse applications, including image classification, semantic segmentation, and object detection, to name a few. For instance, the model consisting of the pre-trained encoder [62, 30] and task-specific decoder is fine-tuned, achieving remarkable performance gain when compared to the model trained from scratch. However, training large, complex models separately for each task and dataset is highly inefficient. Recently, parameter efficient fine-tuning (PEFT) approaches [67, 51, 34] that maximize the efficiency in terms of training parameters have emerged as an alternative to the above-mentioned full fine tuning (FFT) methodologies, achieving competitive performance even with limited computing resources while simplifying the training processes and deployment.
@@
This remarkable progress in vision tasks is primarily driven by approaches including prompt tuning [43] and adapter [33], which have been successfully applied to natural language processing (NLP) tasks. Visual Prompt Tuning (VPT) [35] is the first study to explore the potential of prompt tuning in visual recognition tasks, laying the foundation for the prompt tuning in the field of computer vision. In addition, the adapter-based PEFT methods [10, 65] achieve significant training efficiency by applying the adapter to the Vision Transformer (ViT) and its variants [22, 47, 48].
@@
While most PEFT-based approaches yield performance comparable to baseline methods using the FFT in the image classification task, they do not yet provide sufficiently satisfactory performance to compete with the FFT in other complex downstream tasks. Furthermore, the scalability of prompt-based methods [35, 3] is significantly limited, leading to considerable performance degradation as the number of learnable parameters (i.e., prompts) increases, as reported in [10]. In contrast, adapter-based approaches [10, 34, 65] incorporate lightweight modules to reduce the number of trainable parameters, maintaining stable performance on a range of trainable parameter scales. However, the adapters always apply the same transformation to input features, ignoring individual characteristics of input instances. This may not be an issue when fully tuning whole networks, but it could be a limiting factor in improving performance in adapter-based PEFT methods. Namely, the inflexibility with respect to input instances exacerbates the transfer capability of adapter-based models with only small learnable parts to downstream tasks, limiting their ability to capture unique and task-specific information.
@@
Furthermore, the ViT [22] used in adapter-based models tends to focus on global information rather than fine local details within an image. While this limitation can be partially addressed by employing the Swin Transformer [47], which utilizes local attention mechanisms, the constraint on the number of learnable parameters in adapter-based approaches still restricts the Swin Transformer’s capability to effectively capture local features (Figure 5). Consequently, this negatively affects the performance in dense prediction tasks that require local details.
@@
To address the aforementioned issues, we propose a novel PEFT approach, input-Conditioned transFormer (iConFormer), which leverages a dynamic adapter where parameters are adjusted at the input instance level, unlike existing adapter-based approaches [10, 34, 65]. We introduce an input-Conditioned Network (iCoN) that dynamically generates the parameters for each input feature in the adapter. This approach enables for effectively capturing task-specific and local details for each instance while keeping the number of learnable parameters small. The effectiveness of our method is evidenced by the quantitative analysis in Figure 1. Our method achieves performance competitive to the FFT in both classification and dense prediction tasks including monocular depth estimation, semantic segmentation, and instance segmentation with only additional 1.6% to 2.8% backbone parameters. Our method even surpasses FFT for the image classification in CIFAR100 [38] and the challenging instance segmentation task on COCO [45]. Additionally, iConFormer also outperforms conventional PEFT methods for monocular depth estimation task on NYU-v2 [55], demonstrating the effective utilization of pre-trained backbone parameters with additional learnable parameters dynamically finetuned for specific tasks. We also analyze the capability to capture fine-grained details by visualizing attention maps in Figure 5.
In summary, our contributions are threefold:
@@
Figure 2: Comparison of Full Fine-Tuning (FFT) and the proposed Parameter Efficient Fine-Tuning (PEFT) using iConFormer. (a) FFT, where all parameters are updated during training. (b) Our PEFT (iConFormer), where an dynamic adapter is attached sequentially after the MLP layer in the Transformer. Inside the dynamic adapter, an Input-Conditioned Network (iCoN) generates input-conditioned convolutional kernels in a channel basis, which is detailed in Figure 4. By convolving features with these kernels, iCoN adaptively refines them in accordance with the specific properties of the input, thereby enhancing the model’s capability to effectively process diverse input data in the downstream tasks.
2 Related Work
2.1 Transformer in Vision
@@
Transformers, initially designed for Natural Language Processing (NLP) tasks such as machine translation [56] and text generation [20], have achieved significant success in these areas. This success has led to a shift towards computer vision, starting with the Vision Transformer (ViT) [22]. Subsequently, various Transformer-based models [47, 60, 6, 40, 41, 52, 32] have achieved notable advancements in tasks including image classification [39], semantic segmentation [49, 9, 13, 37], object detection [24, 54, 53], image restoration [21, 46], and depth estimation [25, 14]. Furthermore, transformers have significantly advanced vision recognition through large-scale pretraining [12, 7, 30]. However, their larger size compared to previously prevalent CNN backbones presents challenges for fine-tuning on specific tasks. In this context, our work explores methods to adapt pre-trained transformers into target tasks in a more effective and efficient way.
2.2 Parameter Efficient Fine Tuning
@@
Parameter Efficient Fine-Tuning (PEFT) methods enable the adaptation of large pre-trained models [63, 29, 64, 17, 18] to specific tasks without the need to train the entire model. In NLP, notable approaches include adapter methods [33], which integrate small learnable modules into the model while keeping the pre-trained parameters frozen, with only the added modules being fine-tuned. Additionally, other methods involve tuning specific components such as bias or normalization layers [67, 51], utilizing learnable prompt tokens [43], or applying low-rank matrix approximations [34] to efficiently update parameters. Recently, a method has been proposed to improve inference time while maintaining the training parameter efficiency by selectively skipping the computation of less important tokens [42].
@@
In computer vision, PEFT techniques inspired by NLP have shown significant progress. VPT [35] is the first method to apply prompt tuning approaches to visual recognition tasks. AdaptFormer [10] introduces a parallel adapter framework to enhance the effectiveness of parameter efficient fine-tuning for visual recognition. KAdaptation [31] optimizes the adapter using Kronecker products, and SPT [27] selectively allocates trainable parameters to important positions under a specified budget. In addition, LoRand [65] employs multi-branch low-rank adapters to achieve impressive performance on dense prediction tasks. Our approach is also based on the adapter framework but introduces input-conditioned kernels for instance-specific adaptation, allowing more precise and flexible fine-tuning with a limited number of learnable parameters.
3 Preliminary
3.1 Vision Transformer and its Variants
@@
Vision Transformer (ViT) [22], modified from the Transformer [56] proposed in NLP, integrates image patches and positional encodings to capture spatial information. It consists of a patch embedding layer and multiple sequential encoder blocks, as depicted in Figure 2 (a). Given a batch of images $x\in\mathbb{R}^{B\times H\times W\times 3}$, the patch embedding layer transforms $x$ into sequential patches $x_{p}\in\mathbb{R}^{B\times M\times(P^{2}C)}$, where $H\times W$ is an image resolution, and $B$ is a batch size. $P\times P$ is the resolution of an image patch, $C$ is the output channel, and $M=HW/P^{2}$ is the number of image patches. The patches are linearly embedded into $D$ dimensions to generate the final input $x_{in}\in\mathbb{R}^{B\times M\times D}$.
@@
In the Transformer encoder block, $x_{in}$ is first normalized using LayerNorm (LN) [2], and then processed by a multi-head self-attention layer (MHSA). The output is combined with $x_{in}$ via a residual connection:
@@
This process is repeated $N$ times in the encoder block. In ViT, the self-attention mechanism captures global features by evaluating relationships between all image patches, enabling a comprehensive understanding of complex dependencies. Advancing this approach, Swin Transformer [47, 48] introduces hierarchical attention with shifted windows, which enhances both computational efficiency and feature representation. Other variants [57, 60, 44, 58] leverage multi-scale feature extraction for specific vision tasks, improving the adaptability of Transformer models.
3.2 PEFT Methods
@@
Parameter efficient fine-tuning (PEFT) methods, such as prompt tuning [43, 35], low-rank adaptation [34], and adapters [33, 10], are designed to reduce the number of trainable parameters needed for fine-tuning large models. Here, we briefly review adapter-based PEFT methods related to our approach, which will be detailed in the following section. The adapter introduces small, trainable modules between the layers of the pre-trained model. It can be integrated in sequential or parallel configurations as shown in Figure 3. For instance, if the original architecture processes the features as described in (2), conventional adapters modify this transformation to
$x_{out}=\begin{cases}\gamma\cdot\text{Up}(\sigma(\text{Down}(\tilde{x})))+\acute{x_{in}},&\text{(Sequential)}\\ \tilde{x}+\gamma\cdot\text{Up}(\sigma(\text{Down}(\acute{x_{in}})))+\acute{x_{in}},&\quad\text{(Parallel)}\\ \end{cases}$ (3)
where $\text{Down}(\cdot)$ represents the down-projection of the input features, $\text{Up}(\cdot)$ indicates the up-projection back to the original space, and $\sigma$ denotes an activation function. Here, $\gamma$ is a weighting factor that adjusts the contribution of the adapter output. These approaches allow for task adaptation while minimizing the number of learnable parameters that need to be updated during model training.
4 Proposed Method
4.1 Motivation and Overview
Conventional adapter-based methods [33, 10] rely on static parameter-tuning, where the learnable modules such as ‘Up’ and ‘Down’ added to the original transformation are always static with respect to the input feature $\tilde{x}$ or $\acute{x_{in}}$, as described in (3). Thus, the same transformation is applied to all instances regardless of instance-specific characteristics, limiting the ability to adapt to input feature distributions in the constrained training environment where only a few number of parameters are tuned for various downstream tasks. Namely, in the PEFT that allows for updating only a minimal set of parameters, the static transformations become insufficient for handling diverse input variations, necessitating an instance-level mechanism that dynamically adjusts the learnable parameters conditioned on the input features.
@@ 4.1 Motivation and Overview
To this end, we incorporate dynamic kernel generation, enabling instance-aware transformations while maintaining parameter efficiency under PEFT constraints. Specifically, we introduce iConFormer, a novel framework incorporating the input-Conditioned Network (iCoN), as illustrated in Figure 2 (b). Unlike conventional static adapters, iConFormer dynamically generates input-conditioned convolution kernels tailored to the unique characteristics of each instance, enabling more precise and flexible adaptation. By introducing dynamic kernel generation, the proposed method enhances feature extraction while incorporating locality inductive biases into the pretrained Transformer encoder, allowing the model to dynamically capture both global and local features.
Figure 4: Architecture of input-Conditioned Network (iCoN). The down-projected feature map $\hat{x}$ is used to dynamically generate channel-wise convolution kernels through the iCoN.
4.2 Input-Conditioned Network (iCoN)